OpenStudy (anonymous):

The mass of a child is 35kg. The child is raised to point A which is 1.5m and then released. She swings downwards through the equilibrium position B at 0.80m. What is the loss in gravitational potential energy of the child between A and B. And assuming that air resistance is neglible, calculate the speed of the child as she passes through the equilibrium position B.

OpenStudy (jamesj):

At the top of the swing the child has gravitational potential energy, PE. At the bottom, she has all kinetic energy KE. That KE comes from the PE. Hence KE(bottom) = PE(top). Hence to solve this problem 1. Find PE(top) 2. Set it equal to KE(bottom) 3. Use the formula for KE to find speed, v

OpenStudy (jamesj):

Make sense? Stuck anywhere?

OpenStudy (anonymous):

To find the loss in PE would i use PE = Mgh?

OpenStudy (jamesj):

Yes, PE is calculated as PE = mgh.

OpenStudy (jamesj):

And what's the formula for KE?

OpenStudy (anonymous):

Ok. To calculate The PE = 35*9.8*0.70?

OpenStudy (jamesj):

Why is h = 0.7 m?

OpenStudy (jamesj):

Never mind, I see it now in the problem: it's the difference 1.5 m - 0.8 m

OpenStudy (anonymous):

That,s correct. Am i on the right track?

OpenStudy (jamesj):

Yes

OpenStudy (jamesj):

Hence the PE = 240 Joules

OpenStudy (anonymous):

The second part of the question is that Ek=0.5*mv?

OpenStudy (jamesj):

\[ KE = \frac{1}{2}mv^2 \] yes

OpenStudy (jamesj):

Hence if PE = KE and PE = 240 J, then \[ 240 = \frac{m}{2}v^2 \] or \[ v^2 = 240 \frac{2}{35} = 4202 \]

OpenStudy (jamesj):

*apologies, \[ v^2 = 240\frac{2}{35} = 13.72 \]

OpenStudy (anonymous):

Is 13.72 ms the final answer or would you have to \[\S\] this?

OpenStudy (jamesj):

Absolutely it is not the final answer. You are asked to find the speed of the child, so you must take the square root.

OpenStudy (jamesj):

You need to write down a value for v.

OpenStudy (anonymous):

square root EK/\[half\]m?

OpenStudy (jamesj):

You have a value now for \( v^2 \). We know that \[ v^2 = 13.72 \] So to find \( v \), just take the positive square root of both sides of that equation. Why go backwards in the algebra to KE?

OpenStudy (jamesj):

Hence what is the value of v?

OpenStudy (anonymous):

Hello James. In the value 240 2/35 what is the 2 for?

OpenStudy (jamesj):

We know that the kinetic energy, KE is given by \[ KE = \frac{1}{2}mv^2 \] We know that the KE(bottom) = PE(top). PE(top) = 240 J, hence \[ KE = 240 \] That is \[ \frac{1}{2}mv^2 = 240 \] hence \[ v^2 = \frac{2}{m}240 = \frac{2}{35}240 \] Make sense?

OpenStudy (anonymous):

Yes. One more question if the rope stays taut throughout. Can you explain why the work done by the tension in the rope is zero?

OpenStudy (jamesj):

Wait a second. What's your final value for v?

OpenStudy (anonymous):

3.70ms?

OpenStudy (jamesj):

3.70 m/s, yes. Now I want to show you a quicker way to calculate it.

OpenStudy (anonymous):

Ok carry on!

OpenStudy (jamesj):

We know the PE at the top is equal to the KE at the bottom. As before, let h = vertical height at top of swing v = speed at bottom. Then as PE = KE we have \[ mgh = \frac{1}{2}mv^2 \] So far, so good?

OpenStudy (anonymous):

Yes

OpenStudy (jamesj):

Now we want to solve for v. Notice the mass m cancels from both sides and we have \[ v^2 = 2gh \] Hence \[ v = \sqrt{2gh} \]

OpenStudy (jamesj):

Make sense?

OpenStudy (jamesj):

The point is you now have just one equation to calculate, not two. In your problem h = 0.7 m, hence \[ v = \sqrt{2gh} = \sqrt{2(9.8)(0.7)} = \sqrt{13.72} = 3.70 \] End of problem.

OpenStudy (jamesj):

This formula \[ v = \sqrt{2gh} \] is very handy.

OpenStudy (anonymous):

I know James i have done this equation before and got the same answer but i was'nt sure i was doing it correct though!

OpenStudy (jamesj):

Well now you do. Second part of your question. What is the equation for work done by a force F?

OpenStudy (anonymous):

Yes that,s right!

OpenStudy (jamesj):

I'm addressing now your question: "Can you explain why the work done by the tension in the rope is zero?" So first, what is the definition of Work?

OpenStudy (anonymous):

Work = Pt

OpenStudy (jamesj):

Nope, that's not the definition. The definition is work is the product of Force acting over a distance d, where we only consider the Force in the direction of d. So if I pick up a 1 kg book from the floor and put it on a table 1 m height, the force is F = mg and I acted against a distance of d = 1 meter hence the work I did was Work = Fd = 1(9.8)(1) = 9.8 J

OpenStudy (jamesj):

Now if I pick up the book at hold it one meter off the floor and walk around with it for 8 hours, keeping the book all the time at 1 meter. How much work have I done?

OpenStudy (jamesj):

And I walk 20 kilometers ... how much work have I done on the book?

OpenStudy (jamesj):

still there?

OpenStudy (jamesj):

If you don't know, at least guess. And be quick. Other people are asking for me.

OpenStudy (anonymous):

Sorry james got distracted thanks for your help.